OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 13x - 6y = 20 7x + 4y = 18 2 1 -2 -1

OpenStudy (anonymous):

So you plug in those numbers into the y's in the equation and solve for x, then when you have x solved, plug that answer into the original equation's x's and then solve for y

OpenStudy (abb0t):

Try substitution to solve.

OpenStudy (anonymous):

hey im really tired and this is a pre chapter test so i just need and answer cause im stuck

OpenStudy (abb0t):

Use a term to substitute into equation on the bottom (or top).

OpenStudy (nurali):

13x - 6y = 20----(1) 7x + 4y = 18---- (2) in equation (2) 7x + 4y = 18 7x=18-4y x=18-4y/7 put x in equation (1) 13(18-4y/7)-6y=20 234-52y-42y=20*7 234-94y=140 94y=234-140 94y=94 y=94/94 y=1

OpenStudy (anonymous):

ok tahanks

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

can u help witch one more

OpenStudy (nurali):

i will try.

OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10 2 -1 1 -2

OpenStudy (nurali):

2x + 3y = 1---(1) 3x - 4y = 10---(2) in equation (1) 3y=1-2x y=1-2x/3 put in equation (2) 3x-4(1-2x/3)=10 9x-4+8x=30 17x=30+4 x=34/17 x=2

OpenStudy (anonymous):

thanks i really understand now

OpenStudy (nurali):

My Pleasure.